OpenStudy (anonymous):

If a line with a slope of 4 crosses the y-axis at (0, 1), what is the equation of the line? y − 4x = 1 y + 4x = 1 x − 4y = 1 x + 4y = 1

OpenStudy (sloppycanada):

Equation of a line is y=mx+b. m = slope x and y = (x,y) coordinates b = where the line crosses the y-axis. With this information, do you think you can find the equation of the line?

OpenStudy (anonymous):

no ;(

OpenStudy (dayakar):

can u recognize which is slope and y-intercept in the problem

OpenStudy (anonymous):

no D:

OpenStudy (dayakar):

slope (m) =4 y-intercept (b) =1 which is given in the problem just substitute the above values in y= mx+b u get the answer

OpenStudy (dayakar):

what happens

OpenStudy (anonymous):

x − 4y = 1

OpenStudy (dayakar):

@skyfive 1020 let @Haksorin do some work plz don't give direct answer

OpenStudy (anonymous):

sorry its just that i don't really do mental math more like i look at the problem and instantly know the answer its weird but its just to easy for me i guess